OpenStudy (anonymous):

Generate the first 5 terms of this sequence: f(1) = 2 and f(2) = 3, f(n) = f(n - 1) + f(n - 2), for n > 2. 2, 3, 5, 7, 9 2, 3, 4, 7, 11 2, 3, 5, 9, 11 2, 3, 5, 8, 13

OpenStudy (amoodarya):

f(3)=f(2)+f(1)=2+3=5 f(4)=f(3)+f(2)=5+3=8 f(5)=f(4)+f(3)=8+5=13 can you go on ?

OpenStudy (anonymous):

i dont even know how you got that

OpenStudy (amoodarya):

it is easier to say : every number in this sequence = sum of two term before

OpenStudy (anonymous):

oh ok. thank you

OpenStudy (anonymous):

so the last one?

OpenStudy (amoodarya):

let's do it again : put n=3 in relation f(3)=f(3-1)+f(3-2) --->f(3)=f(2)+f(1) now put f(2) ,f(1)

OpenStudy (anonymous):

oh ok. makes more sense now

OpenStudy (amoodarya):

now , can you tell me the next number is ?